OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify -27/32 * 14/15 * 3

OpenStudy (dumbcow):

multiply straight across cancel when possible ex. 14/32 = 7/16 3/15 = 1/5 -> (-27*7)/(16*5) = -189/80

OpenStudy (anonymous):

thank you so much